an object is thrown from ground with speed 20m/s at an angle 30 degree with the horizontal. its centripetal acceleration one second after projection in SI unit is :
1)10
2)0
3)5
4)12
 

Grade:11

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

Centripetal accel is provided by the gravity
and is equal to gcos theta where theta is angle made by
ball's velocity at that moment.
vsin30-gt=10-10=0
=>at that moment its making angle zero with horizontal
=>gcos0=g

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
manish gupta
9 years ago
an object is thrown from ground with speed 20m/s at an angle 30 degree with the horizontal. its centripetal acceleration one second after projection in SI uni